OpenStudy (anonymous):

investor invested $2,200 in two mutual funds. one fund earned 9% and the other earned 2% profit. If total profit was $184 how much was invested in each fund

OpenStudy (barrycarter):

How far did you get?

OpenStudy (anonymous):

i dont know where to start is there a formula

OpenStudy (barrycarter):

LOL :) You don't need a formula here. Let x=amount invested in first mutual fund, y=in second. You know x+y=2200 and what else?

OpenStudy (anonymous):

am lost

OpenStudy (helder_edwin):

if u invest M dollars and u r paid r% then your profit is \[ \large r\%\cdot M=\frac{r}{100}\cdot M \] use this to form a second equation in addition to the one BARRYCARTER just gave yoiu

OpenStudy (anonymous):

i hate math lol

OpenStudy (helder_edwin):

the first mutual fund pays you 9% then your profit is \[ \large 9\%\cdot x=0.09x \] the second one pays you 2% then your profit is \[ \large 2\%\cdot y=0.02y \] Then \[ \large 0.09x+0.02y= \]

OpenStudy (anonymous):

but how do you find x or y

OpenStudy (helder_edwin):

u have to solve the linear system (two equations) u now have. but before \[ \large 0.09x+0.02y=?? \]

OpenStudy (anonymous):

0.11

OpenStudy (helder_edwin):

no, read the problem. how much is your total profit??

OpenStudy (anonymous):

184

OpenStudy (helder_edwin):

great. now u have the system \[ \large \begin{cases} x+y=2200\\ 0.09x+0.02y=184 \end{cases} \] solve it!

OpenStudy (anonymous):

x=2000 y=200

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

thanks you

OpenStudy (helder_edwin):

u r welcome. don't hate math, it can be fun. it could help u getting a well paying job.

OpenStudy (anonymous):

i know I just get frustrated and not easy being a single mom working full time and trying to attend college :)

OpenStudy (helder_edwin):

the best of luck to you.

OpenStudy (anonymous):

thanks :)